OpenStudy (anonymous):

D E F and A B C are complementary angles and A B C is nine times as large as D E F. Determine the measure of each angle

ganeshie8 (ganeshie8):

81, 9

OpenStudy (anonymous):

Algebraically this is what you have: Let the measurement of the angle DEF be x and the measurement of the angle ABC be y. For x and y to be complimentary means that when you add them together you get exactly 90 degrees (a right angle). So you have the equation x+y=90 . Now, you also have that y is 9 times as large as x, which boils down to having the equation y=9x . So now you have the equations x+y=90 y=9x Two equations and two unknowns perfect so now you can solve for what x and y are. So you simply substitute y=9x into x+y=90 Step 1) x+y=90 Step 2)Substitute y=9x: x+(9x)=90 Step 3) Simplify x+9x=90 10x=90 x=9 Step 4) solve for y, y=9x y=9(9)=81 so x=9 and y=81 I'm assuming these are in degrees.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

you're welcome :-)